# An upper bound of the silhouette validation metric for clustering

### v1

@@ Abstract
The silhouette coefficient summarizes, per observation, cohesion versus separation in [1,1][-1,1]; the average silhouette width ($\mathrm{ASW}$) is a common internal measure of clustering quality where higher values indicate more coveted results. However, the dataset-specific maximum of $\mathrm{ASW}$ is typically unknown, and the standard upper limit $1$ is often unattainable. In this work, we derive for each data point in a given dataset a sharp upper bound on its silhouette width. By aggregating these individual bounds, we present a canonical data-dependent upper bound on $\mathrm{ASW}$ that often assumes values well below $1$. The presented bounds can indicate whether individual data points can ever be well placed, enable early stopping of silhouette-based optimization loops, and help answer a key question: How close is my clustering result to the best possible outcome on this specific data? Across synthetic and real datasets, the bounds are provably near-tight in many cases and offer significant enrichment of cluster quality evaluation.
@@ 1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [21]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [42].
Among cluster validation indices, the silhouette—also known as the silhouette coefficient, silhouette score, or silhouette index—is one of the most widely used measures for assessing how well each point fits within its assigned cluster relative to the nearest neighboring cluster [74]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well-separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [11] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [67] proposed alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics can impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Directly comparing $\mathrm{ASW}$ values between such datasets can therefore lead to misguided conclusions.
@@
A further obstacle is the combinatorial explosion of the partition space. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion, including the silhouette index, computationally prohibitive. [11] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a principled and data-aware benchmark, namely a theoretical bound, against which the quality of such locally optimal solutions can be evaluated.
Figure 1: Illustration of the silhouette coefficient calculation for $x_{1}$, showing intra-cluster distances within cluster $C_{1}$ and inter-cluster distances to cluster $C_{2}$.
In general, the major contributions of this paper are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$ to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass. We also derive for each data point an upper bound on its silhouette width. These per-point bounds have the potential to highlight observations that cannot be well-placed in any clustering. The bounds are computed by an $O(n^{2}\!\log n)$ algorithm that is practical for medium– to large-scale datasets. In some scenarios, when combined with empirical results, the proposed $\mathrm{ASW}$ upper bound can serve as proof that a silhouette-optimal solution has actually been found.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures and clustering baselines. The bound is tight or near-tight in many cases, diagnoses weak clusterability when no inherent group structure exists, tightens under minimum cluster-size constraints, and enables early stopping in silhouette-based model selection.
    All datasets, preprocessing scripts, bound-computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
@@ 2 Related work
Silhouette coefficientEnergy & Power Systems[62][73][93][23][27][91][51][94][40][2][19]Transportation & Mobility[60][86][8][79][37][66][58][56][4]Biomedical & Health[43][78][33][34][9][47][75][44][87][12]Food & Agriculture[25][81][32]Social & Economic[63][72][36][90][5]Materials & Chemistry[69][13][76][6][17]Environment & Earth Sciences[61][38][89][41][10][14][35][39][92][18]Sensor Networks & IoT[49][50][64][65][85]Big Data[70][83][82][16]Cluster Validity & Selection[45][46][11][55][15][30][7]Algorithm Development[48][80][59][71][90][8][79][49][50][22][29][31][28][26][32]Feature & Dim-Reduction[75][24][95][16]
@@
The silhouette coefficient [74] is a distance-based internal validity index that quantifies how well each object fits its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}\in C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$ (its average dissimilarity to members of $C_{\ell}$) and nearest-cluster separation by $b(x_{i})$ (its average dissimilarity to the closest rival cluster); see Sec.  3 for formal definitions. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own.
Aggregating over points gives the average silhouette width (ASW), a scale-free summary of partition quality (Sec.  3). In practice, ASW is often used for internal model selection by choosing the number of clusters $K$ that maximizes it; rule-of-thumb interpretations treat larger ASW as evidence of well-separated, compact clusters, and values near zero or negative as signs of overlap or a poor choice of $K$. Despite numerous extensions—e.g., to categorical data [20], fuzzy clustering [53], hierarchical clustering [63], model-based clustering [35], spectral clustering [87], spherical clustering [18], and optimization-based methods [88]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
@@ 3 Preliminaries
Table 1: A toy dataset of five samples with two features.
Point $Att_{1}$ $Att_{2}$
$x_{1}$ 1.0 2.0
$x_{2}$ 2.0 1.0
$x_{3}$ 1.5 2.5
$x_{4}$ 6.0 2.0
$x_{5}$ 6.0 3.0
Figure 3: Scatter plot of the toy dataset.
We consider a dataset of $n$ observations described by $m$ features. Throughout, we fix an integer $n\geq 4$ and index the observations by the set $\{1,\dots,n\}$. For example, Tab. 3 and Fig. 3 present a dataset consisting of five data points, each described by two features. We use this toy dataset for illustrative examples in this section.
When applicable, the samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$ of a dataset can be stored in a matrix of the form
$X\;=\;[x_{ij}]_{1\leq i\leq n,\;1\leq j\leq m}\;\in\;\mathbb{R}^{n\times m},\qquad\mathbf{x}_{i}:=(x_{i1},\dots,x_{im})^{\top}\in\mathbb{R}^{m}.$
We deal with datasets where there is a measure of dissimilarity (or distance) between each pair of data points. All pairwise dissimilarities are contained in the datasets dissimilarity matrix.
@@
    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non-negative),
@@
The triangle inequality is not required. We additionally exclude matrices containing a row of all zeros.
@@
For instance, if the Euclidean distance is used, then $\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2}$. The dissimilarity matrix corresponding to the dataset shown in Fig. 3 is:
$\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2},\qquad\Delta=\begin{pmatrix}0&1.414&0.707&5.000&5.099\\ 1.414&0&1.581&4.123&4.472\\ 0.707&1.581&0&4.528&4.528\\ 5.000&4.123&4.528&0&1.000\\ 5.099&4.472&4.528&1.000&0\end{pmatrix}$
We remark that the upper bound proposed in Sec. 4 depends only on the dissimilarity matrix of the dataset; no observation vectors are required. Hence, it is just as data-agnostic as the $\mathrm{ASW}$ itself.
@@
The sorted off-diagonal dissimilarity matrix for the toy dataset is shown in Tab. 2.
Table 2: Row-wise sorted off-diagonal distances $\hat{\Delta}$ corresponding to the toy dataset.
Point $\hat{\Delta}_{i1}$ $\hat{\Delta}_{i2}$ $\hat{\Delta}_{i3}$ $\hat{\Delta}_{i4}$
$x_{1}$ 0.707 1.414 5.000 5.099
$x_{2}$ 1.414 1.581 4.123 4.472
$x_{3}$ 0.707 1.581 4.528 4.528
$x_{4}$ 1.000 4.123 4.528 5.000
$x_{5}$ 1.000 4.472 4.528 5.099
@@
A popular clustering method is the K-means algorithm, which we will use in Sec. 5. Its objective is to minimize the within-cluster sum of squares (or variance).
Definition 3 ($K$-means [57]).
When the desired number of clusters $K$ is specified, the classical $\mathbf{K}$-means problem seeks
$\operatorname*{argmin}_{\mathcal{C}_{K}}\sum_{I=1}^{K}\;\sum_{i\in C_{I}}\bigl{\lVert}\mathbf{x}_{i}-\boldsymbol{\mu}_{I}\bigr{\rVert}_{2}^{2},\qquad\boldsymbol{\mu}_{I}\coloneqq\frac{1}{|C_{I}|}\sum_{i\in C_{I}}\mathbf{x}_{i}.$
Figure 5: Clustering results using $K$-means on the dataset in Fig. 3.
For instance, Fig. 5 illustrates the result of applying $K$-means clustering to the toy dataset with $K=2$ clusters.
Alternative objectives such as linkage-based agglomerative clustering may be formulated directly in terms of $\Delta$, but the underlying aim remains to form internally cohesive and externally separated groups. For instance, Fig.6 shows a dendrogram obtained from performing hierarchical clustering on the dissimilarity matrix corresponding to the toy dataset.
Figure 6: Hierarchical clustering on the distance matrix of data points in Fig. 3.
We will now define the silhouette, which was introduced in 1987 by P.J. Rousseeuw [74]. We employ similar notation.
@@
To illustrate with an example, the silhouette value of each data point from the clustering results shown in Fig. 5 is displayed in Tab. 3 and Fig. 7. The corresponding $\operatorname{ASW}$ is $0.7512$, indicating well-separated and internally cohesive clusters.
Table 3: Silhouette components for each point in the toy dataset after applying K-means clustering with $K=2$.
$i$ $a(i)$ $b(i)$ $s(i)$
$x_{1}$ 1.061 5.050 0.790
$x_{2}$ 1.498 4.298 0.652
$x_{3}$ 1.144 4.528 0.747
$x_{4}$ 1.000 4.550 0.780
$x_{5}$ 1.000 4.700 0.787
Figure 7: Silhouette plots of clustering results from Fig. 5.
Individual silhouette widths provide a valuable graphical diagnostic (see Fig. 7), but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [54]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, the limitation often stems from the intrinsic structure of the dissimilarity data rather than a suboptimal choice of algorithm. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
@@
Once again, recall the toy data (Tab. 3) and its dissimilarity matrix $\Delta$. We evaluate $\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ for every admissible clustering with $K\in\{2,3,4,5\}$ (all set partitions except the trivial K=1K{=}1 case). Among the 51 candidates  [^3] The fifth Bell number $B_{5}=52$ and excluding K=1K{=}1 gives $51$., the maximizer is unique:
$\mathcal{C}^{\ast}\;=\;\bigl{\{}\{x_{1},x_{2},x_{3}\},\;\{x_{4},x_{5}\}\bigr{\}}\quad\text{with}\quad\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)\;=\;0.7512,$
which matches the per-point silhouettes reported in Tab. 3 and Fig. 7. Finally, the best achievable $\operatorname{ASW}$ within each fixed $K$ is depicted in Tab. 4.
Table 4: Optimal $\operatorname{ASW}$ within each $K$ on the toy dataset.
$K$ Best partition Best ASW
2 $\{x_{1},x_{2},x_{3}\}\mid\{x_{4},x_{5}\}$ 0.7512
3 $\{x_{1},x_{3}\}\mid\{x_{2}\}\mid\{x_{4},x_{5}\}$ 0.5173
4 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4},x_{5}\}$ 0.3068
5 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4}\}\mid\{x_{5}\}$ 0.0000
4 The proposed upper bound of the silhouette
4.1 Overview
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and moreover, there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and supports optional minimum-cluster-size constraints to make it even sharper.
4.2 Derivation

### v2

@@ Abstract
The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in [1,1][-1,1]. The average silhouette width (ASW) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters. However, the dataset-specific maximum of ASW is typically unknown, and the standard upper limit of 1 is rarely attainable. In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound on the ASW. This bound—often substantially below 1—enhances the interpretability of empirical ASW values by indicating how close a given clustering result is to the best possible outcome on that dataset. It can be used to confirm global optimality, guide the evaluation of clustering solutions, and be refined to incorporate minimum cluster-size constraints for greater practical relevance. Finally, we extend the framework to establish a corresponding bound for the macro-averaged silhouette.
@@ 1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [aggarwal2013introduction]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [dinh2025data]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [jain1999data].
Among the cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures to assess how well each point fits within its assigned cluster relative to its nearest neighboring cluster [ogsilhouette]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well‑separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [batool2021clustering] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [john] examined alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Comparing $\mathrm{ASW}$ values across datasets with different properties can therefore lead to misguided conclusions about cluster quality.
@@
The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive. [batool2021clustering] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated.
Figure 1: Illustration of the silhouette coefficient calculation.
The main contributions are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures. We show that the bound is nearly tight in many of the cases examined.
    All datasets, preprocessing scripts, bound computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
@@ 2 Related work
Silhouette coefficientEnergy & Power Systems[ouyang2024k][rodriguez2024novel][zheng2025practical][duarte2021increasing][eskandarnia2022embedded][zhao2023generalized][koutsandreas2024harnessing][zhou2025recognition][hu2023transformer][alrasheed2024building][diaz2022typologies]Transportation & Mobility[mohd2024passengers][wang2021analyzing][bai2023adaptive][sheng2025multi][guo2025adaptive][passarella2024anomaly][ma2025unravelling][li2025clustering][ansari2021spatiotemporal]Biomedical & Health[jamalpour2025high][sharafoddini2021identifying][ghaderi2023self][gill2024symptom][baier2024nocturnal][kim2022pathological][rustam2022data][jha2021apache][xian2025systematic][bhagat2022dpcf]Food & Agriculture[edris2024identifying][singh2025development][gere2023recommendations]Social & Economic[owoo2025hierarchical][quintero2022analysis][gulma2022new][zhang2025two][aonso2024cluster]Materials & Chemistry[pham2025clustering][bi2025comprehensive][sancho2022cluster][bagirov2023novel][consuegra101unsupervised]Environment & Earth Sciences[nourani2022robust][guo2023combining][yang2022zoning][huang2023dominant][balbi2024hierarchical][biabiany2020design][guan2025novel][gupta2025criteria][zhao2023comparison][deng2025on]Sensor Networks & IoT[kotary2020distributed][kotary2021many][panwar2024distributed][panwar2025distributed][waidyanatha2023regularity]Big Data[qaroush2025maximum][thirumoorthy2021hybrid][song2023novel][chander2023dolphin]Cluster Validity & Selection[khan2025standardizing][khan2025standardization][batool2021clustering][lenssen2024medoid][capo2023fast][gagolewski2021are][bagirov2023finding]Algorithm Development[koren2025preprocessing][silva2025automatic][manochandar2020development][qin2025single][zhang2025two][bai2023adaptive][sheng2025multi][kotary2020distributed][kotary2021many][du2024multi][fu2025gsscl][gan2024discrimination][fu2023gc][enriquez2025performance][gere2023recommendations]Feature & Dim‑Reduction[rustam2022data][dwivedi2025novel][zhou2024ocae][chander2023dolphin]
@@
The silhouette coefficient [ogsilhouette] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}$ in cluster $C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{\ell}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own. Aggregating over points gives the ASW, a dataset-level summary of partition quality.
In practice, ASW is often used for model selection by choosing the number of clusters $K$ that maximizes it. Despite numerous extensions—e.g., to categorical data [dinh2019estimating], fuzzy clustering [lasek2024employing], hierarchical clustering [owoo2025hierarchical], model-based clustering [guan2025novel], spectral clustering [xian2025systematic], spherical clustering [deng2025on], and optimization-based methods [yang2024clustering]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
Other commonly used internal validation indices include the Davies–Bouldin index [dbindex] and the Calinski–Harabasz index [chindex], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains. To account for cluster-size imbalance, we additionally employ the macro-averaged silhouette, defined as the unweighted mean of per-cluster silhouette averages. This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
Silhouette-optimization
PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [kaufman90], designed to directly optimize the ASW [pamsil]. Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize ASW. It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in ASW and terminates once no further improvement is possible, ensuring convergence to a local optimum. An efficient implementation of PAMSIL is available in the open-source K-Medoids Python package [Schubert2022]. Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local ASW optima [batool2021clustering].
External validation indices
When ground-truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition. Two of the most commonly used indices are the adjusted Rand index (ARI) [ari] and the adjusted mutual information (AMI) [ami]. Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
The ARI extends the classical Rand index by comparing all pairs of samples and counting how consistently they are assigned to the same or different clusters in both partitions. Its normalization accounts for random overlap, producing values in [1,1][-1,1] with $1$ indicating identical partitions. The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
Both indices are symmetric, easy to interpret, and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
@@ 3 Preliminaries
Let $n\in\mathbb{N}$ with $n\geq 4$. We consider a dataset consisting of $n$ objects indexed by the set $\{1,\ldots,n\}$. Each pair of objects is associated with a measure of dissimilarity (or distance), and all pairwise dissimilarities are collected in the dataset’s dissimilarity matrix.
@@
    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non‑negative),
@@
The triangle inequality is not required. We also exclude matrices containing a row of all zeros.
@@
We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [ogsilhouette]. We employ similar notation.
@@
Individual silhouette widths provide a valuable graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [medoid]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
@@
4 A data-dependent upper bound of the silhouette
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.